OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope. (8, 3); m = 6

OpenStudy (anonymous):

do you know what the point-slope formula is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the point slope formula is: y-y1 = m(x-x1) where (x1, y1) is a point that the line passes through, and m is the slope of the line plug in what you know and youve got an answer

OpenStudy (anonymous):

y + 3 = 6x + 8 wout hat b it ?

OpenStudy (anonymous):

you forgot the parenthesis around (x+8) also, the formula says to subtract. you added the correct equation is y-3=6(x-8)

OpenStudy (anonymous):

o man thanks !!! imma study this more